integrate sqrt(x ^ 3 + 2x ^ 2 + x) dx


1
Expert's answer
2021-06-15T10:17:02-0400


\int\sqrt{(x^3+2x^2+x)}dx\newline


\int\sqrt{x(x^2+2x+1)}dx\newline


\int\sqrt{x(x+1)(x+1)}dx\newline


\int\sqrt{x(x+1)^2}dx\newline


\int\sqrt{x}(x+1)dx\newline



\int x\sqrt{x}dx + \int\sqrt{x}dx\newline



\int x^{\frac{3}{2}}dx + \int x^\frac{1}{2}dx\newline



\frac{x^{\frac{3}{2}+1}}{\frac{3}{2}+1} + \frac{x^{\frac{1}{2}+1}}{\frac{1}{2}+1}+c\newline


\frac{2x^{\frac{5}{2}}}{5} + \frac{2x^{\frac{3}{2}}}{3}+c\newline